# Impact of Exchange-Correlation Functionals on Predictions of Phonon Hydrodynamics: A Study of Fluorides, Chlorides, and Hydrides

### v2

@@ 3.1 General Properties
$Z^{*}$ 0.99 0.98 0.97 1.004 (GGA)ba, 0.975 (LDA)ba
$\omega_{\mathrm{LO}}$ 22.78 23.35 24.31
$\omega_{\mathrm{TO}}$ 12.93 14.14 15.67 14.8 (CN)az, 12.6 (QH)az
$E_{g}$ 3.43 3.25 3.04 5.85 (GW)bc, 3.18 (LDA)bc
aRef. [wang2012phase]. bRef. [hartel2010relaxation]. cRef. [prencipe1995ab]. dRef. [solovyeva2016alchemical]. eRef. [messaoudi2015band]. fRef. [pies2006crystal]. gRef. [bruno2014crc]. hRef. [lowndes1969dielectric]. iRef. [shukla2000ab]. jRef. [born1996dynamical]. kRef. [buyers1967lattice]. lRef. [wu2017electronic]. mRef. [brown1970extreme]. nRef. [lindsay2016isotope]. oRef. [liang2018lattice]. pRef. [hou2016structural]. qRef. [ekinci2004thermal]. rRef. [wyckoff1966crystal]. sRef. [bernardini1998electronic]. tRef. [dolling1968lattice]. uRef. [guo2008ab]. vRef. [shirley1996detailed]. wRef. [roessler1967electronic]. xRef. [piacentini1976thermoreflectance]. yRef. [buhrer1970lattice]. zRef. [hass1963infrared]. aaRef. [cunningham1974second]. abRef. [pantelides1975electronic]. acRef. [froyen1986structural]. adRef. [paier2006screened]. aeRef. [heyd2004efficient]. afRef. [togo2022mode]. agRef. [raunio1969phonon]. ahRef. [miyata1968optical]. aiRef. [himpsel1978angle]. ajRef. [copley1969lattice]. akRef. [erdinc2015ab]. alRef. [van2007electronic]. amRef. [yang2017thermal]. anRef. [verble1968lattice]. aoRef. [pretzel1960properties]. apRef. [messer1960survey]. aqRef. [anderson1970isotopic]. arRef. [blat1991calculations]. asRef. [brodsky1967infrared]. atRef. [laplaze1976etude]. auRef. [roma1996phonon]. avRef. [kondo1988effect]. awRef. [plekhanov1998wannier]. axRef. [ke2005decomposition]. ayRef. [sun2012structural]. azRef. [martins1990equations]. baRef. [barrera2005lda]. bbRef. [humphries2016fluoride]. bcRef. [lebegue2003implementation].
@@ 3.2 Mechanical Properties
The mechanical properties of fluorides, chlorides and hydrides, including the bulk modulus and elastic constants, are summarized in Table 2. For cubic crystals, the three independent elastic constants are $C_{11}$, $C_{12}$, and $C_{44}$. Our calculated bulk modulus and elastic constant values are in good agreement with the experimental values. We computed the bulk modulus $B$ and the elastic constants $C_{ij}$ at $0\,\mathrm{K}$. LDA typically yields the largest bulk moduli, PBESOL is intermediate, and PBE gives the smallest (e.g., LiF: $B_{\mathrm{PBE}}=67.17\,\mathrm{GPa}$, $B_{\mathrm{PBESOL}}=72.55\,\mathrm{GPa}$, $B_{\mathrm{LDA}}=86.78\,\mathrm{GPa}$) . This ordering follows from their equilibrium volumes: LDA overbinds (smaller lattice constants), which increases the curvature of the energy–volume relation $B=V\,\frac{\mathrm{d}^{2}E}{\mathrm{d}V^{2}}$ [rodney2005discrete]; PBE underbinds (larger volumes), which reduces that curvature, and PBESOL was tuned for solids and therefore falls between. Replacing Li with Na and then K systematically softens the lattice at a fixed anion. For fluorides, using PBE for comparisons, the bulk modulus decreases from $B=67.17\,\mathrm{GPa}$ (LiF) to $46.79\,\mathrm{GPa}$ (NaF) and $29.01\,\mathrm{GPa}$ (KF). For chlorides, it decreases from $23.67\,\mathrm{GPa}$ (NaCl) to $16.27\,\mathrm{GPa}$ (KCl). For hydrides, it decreases from $34.24\,\mathrm{GPa}$ (LiH) to $22.16\,\mathrm{GPa}$ (NaH) and $13.74\,\mathrm{GPa}$ (KH). In magnitude, fluorides are the stiffest for a given cation, chlorides are intermediate, and hydrides are the softest for Na and K.
@@
For cubic crystals, mechanical stability requires the elastic constants to satisfy four conditions: $C_{11}>0$, $C_{44}>0$, $C_{11}-C_{12}>0$, and $C_{11}+2C_{12}>0$ [hou2016structural]. The calculated values of the elastic constants for the fluoride, chloride, and hydride materials satisfy all these conditions. Comparing theoretical elastic constants with experiments is challenging because of the uncertainties in experimentally determining $C_{12}$, which is a linear combination of the elastic constants in the <110> direction, whereas $C_{11}$ and $C_{44}$ values can be measured directly by techniques like ultrasonics [cohen1975theory]. Directional elasticity is captured by the Zener anisotropy factor $AF=2C_{44}/(C_{11}-C_{12})$, where $AF=1$ for an elastic isotropic medium [zener1949elasticity]. Using PBE elastic constants, the Zener anisotropy factors are: LiF $AF\!\approx\!1.90$, LiH $AF\!\approx\!1.65$, and NaH $AF\!\approx\!1.45$ are anisotropic; KH $AF\!\approx\!0.94$ is nearly isotropic; NaF $AF\!\approx\!0.72$, NaCl $AF\!\approx\!0.69$, KF $AF\!\approx\!0.61$, and KCl $AF\!\approx\!0.44$ are isotropic. This matches the $AF$ values reported by Sirdeshmukh et al. [sirdeshmukh2001alkali].
@@
From the elastic constants, one can also inspect the data to determine whether the material is brittle or ductile. According to Cauchy pressure ($C_{12}$$C_{44}$), if the difference is positive, the material is ductile; otherwise, it is brittle [pettifor1991bonding]. Based on this, all eight materials are brittle (i.e., negative difference).
@@
Table 2: Bulk modulus $B$ (GPa) and elastic constants $C_{ij}$ (GPa) of calculated (PBE, PBESOL, LDA), other theoretical data, and experimental data for NaF, LiF, KF, NaCl, KCl, LiH, NaH and KH.
@@
$B$ 67.17 72.55 86.78 87.1125c, 75.9e, 67.71 (GGA)j, 73.59 (LDA)j, 72.99 (Birch–Murnaghan EOS)k, 70.53 (Vinet EOS)k 66.4f, 76.9g
$C_{11}$ 109.45 126.71 143.12 125.8e, 224.83i, 153.15k 113.73f, 135.8g
$C_{12}$ 46.03 45.48 50.61 50.9e, 121.09i, 45.62k 47.59f, 47.4g
$C_{44}$ 60.32 61.91 68.07 76.1e, 121.09i, 55.08k 63.68f, 68.7g
@@
$B$ 13.74 14.67 17.40 13.4 (GGA)p, 17.4 (LDA)p, 13.3q, 19.0r 16.0r, 15.6 $\pm$ 1.5aa
$C_{11}$ 28.32 31.81 39.50 39.4p, 26.8q
$C_{12}$ 6.45 6.11 6.35 6.3p, 6.5q
$C_{44}$ 10.32 10.09 10.43 10.4p, 10.6q
aRef. [wang2012phase]. bRef. [solovyeva2016alchemical]. cRef. [otero2011gibbs2]. dRef. [recio1990theoretical]. eRef. [prencipe1995ab]. fRef. [miller1964pressure]. gRef. [haussuhl1960thermo] (data taken from Ref. [prencipe1995ab]). hRef. [lewis1967elastic] (bulk modulus values from Ref. [solovyeva2016alchemical], measured at $T=4.2$ K). iRef. [sangster1978interionic]. jRef. [liang2018lattice]. kRef. [hou2016structural]. lRef. [froyen1986structural]. mRef. [copley1969lattice]. nRef. [zhang2007first]. oRef. [bouhadda2007first]. pRef. [barrera2005lda]. qRef. [zhang2007phonon]. rRef. [rodriguez1992pressure]. sRef. [gerlich1974pressure]. tRef. [pandey1985intrinsic]. uRef. [haque1990hydrides]. vRef. [dyck1981lattice]. wRef. [sun2011ab]. xRef. [sun2012structural]. yRef. [ojwang2008modeling]. zRef. [duclos1987high]. aaRef. [hochheimer1985high].
@@ 3.3 Thermal Properties
We calculate the room temperature lattice thermal conductivity, specific heat, average Grüneisen parameter, and density values at room temperature. The computed values are listed in Table 3 and compared with theoretical and experimental values from the literature. PBESOL generally yields thermal conductivities $k$ that lie between those predicted by PBE (lower) and LDA (higher). Some examples include: for NaF, PBE (18.5 W.m-1.K-1) matches experiment (18.4, 18.7) [morelli2006high, smirnov1845thermal], while LDA overestimates (31.5); for KCl, PBESOL (6.90 W.m-1.K-1) tracks experiments (7.1, 6.7) [morelli2006high, sirdeshmukh2001alkali]; for KF and NaCl, LDA (7.32 and 7.86 W.m-1.K-1) is the closest to the measured values (7.1) [morelli2006high, sirdeshmukh2001alkali]. Specific heats are generally in good agreement with the expeirement values across functionals (e.g., for NaCl $c_{p}\approx 0.82\text{--}0.83~\mathrm{kJ\,.kg^{-1}\,.K^{-1}}$ vs. $0.86~\mathrm{kJ\,.kg^{-1}\,.K^{-1}}$ [haynes2016crc]). Average Grüneisen parameters fall within 1.4–2.1 for all materials. On the other hand, density orders across the series (fluorides highest, chlorides intermediate, hydrides lowest), and the computed values closely track the experimental results (PBESOL performs best for halides, while PBE performs best for hydrides). For instance, LiF, PBESOL predicts a value of 2.67 g.cm-3, closely matching experiment (2.639) [miller1964pressure], whereas KH, PBE predicts a value of 1.44 g.cm-3, closely matching experiments (1.43) [haynes2016crc, mueller2013metal].
@@
The lattice thermal conductivity is calculated for temperatures ranging from 100 K to 1000 K (see Fig.  3) and compared with the available data from the literature. The effect of isotopes on the thermal conductivity calculations is also shown in the same figure. All compounds show the expected monotonic decrease of $k$ with $T$. Including naturally abundant isotope scattering in the calculations mainly depresses $k$ (below 200 K) and has less impact at high $T$, where the three functionals produce nearly parallel slopes. Isotope scattering can significantly impact the lattice thermal conductivity $k$ in materials with mixed natural isotopes, but has little to no effect when the constituent elements are effectively monoisotopic. For example, NaF shows essentially no isotope effect because the natural abundance of Na is $\sim 100\%$ $^{23}\mathrm{Na}$ and the natural abundance of F is $\sim 100\%$ $^{19}\mathrm{F}$, whereas LiF exhibits a reduction in $k$ when isotopes are included due to its mixed lithium composition ($\sim 92.5\%$ $^{7}\mathrm{Li}$, $\sim 7.5\%$ $^{6}\mathrm{Li}$ [tomascak2004developments]) that introduces phonon mass-disorder scattering most evident at low $T$.
@@
Figure 3: Calculated lattice thermal conductivity for: (a) NaF. (b) LiF. (c) KF. (d) NaCl. (e) KCl. (f) LiH. (g) NaH. (h) KH. Calculations were done with PBE (black solid line), PBESOL (red solid line), and LDA (blue solid line). Calculations also show the effect of the isotopes on the thermal conductivity: PBE (black circles), PBESOL (red circles), and LDA (blue circles). Experimental data: Ref. [smirnov1845thermal] (orange circles), Ref. [petrov1974temperature] (gray triangles), Ref. [walker1963thermal] (pink squares), Ref. [petrov1976behaviour] (orange triangle), Ref. [vetrano1957batelle] (pink triangles), Ref. [slack1973nonmetallic] (orange squares). Theoretical data: Ref. [gheribi2016formulation] (purple diamonds), Ref. [lindsay2016isotope] (pink diamonds), Ref. [liang2018lattice] (green pentagons), Ref. [singh2003effects] (purple squares), Ref. [yang2017thermal] (purple pentagons).
@@
Table 3: Thermal conductivity $k$ ($\mathrm{W\,m^{-1}\,K^{-1}}$), specific heat $c_{p}$ ($\mathrm{kJ\,kg^{-1}\,K^{-1}}$), average Grüneisen parameter $\bar{\gamma}$, and density $\rho$ ($\mathrm{g\,cm^{-3}}$) of calculated (PBE, PBESOL, LDA), other theoretical data, and experimental data for NaF, LiF, KF, NaCl, KCl, LiH, NaH and KH at room temperature.
@@
$k$ 8.88 11.07 13.91 19.4a, 8.71b, 22.42 ($a_{0}=3.886$ Å)j, 14.59 ($a(300\,\mathrm{K})=4.004$ Å)j, 13.89 ($a_{0}=4.002$ Å)k, 13.48 ($a(300\,\mathrm{K})=4.03$ Å)k 17.6a, 14.09l, 14.2m
$c_{p}$ 1.62 1.59 1.55 1.62d 1.54d, 1.62m
$\bar{\gamma}$ 2.08 1.88 1.54 2.02b, 1.48e, 1.24 ($a_{0}=4.002$ Å)k, 1.36 ($a(300\,\mathrm{K})=4.03$ Å)k 1.5a, 1.64e
$\rho$ 2.56 2.67 2.88 2.64f 2.639h, 2.64m
@@
$k$ 4.23 3.31 10.66
$c_{p}$ 0.99 0.98 0.94
$\bar{\gamma}$ 1.59 1.55 1.28 2.0 (CN)x, 2.2 (QH)x
$\rho$ 1.44 1.51 1.63 1.432q 1.43i, 1.43q
aRef. [morelli2006high]. bRef. [toher2014high]. cRef. [smirnov1845thermal]. dRef. [touloukian1972thermophysical]. eRef. [kachhava1968semiempirical]. fRef. [spangenberg1957elastischen]. gRef. [lewis1967elastic]. hRef. [miller1964pressure]. iRef. [haynes2016crc]. jRef. [lindsay2016isotope]. kRef. [liang2018lattice]. lRef. [petrov1976behaviour]. mRef. [sirdeshmukh2001alkali]. nRef. [yang2017thermal]. oRef. [vetrano1957batelle]. pRef. [yates1974specific]. qRef. [mueller2013metal]. rRef. [islam1994crystal]. sRef. [gerlich1974pressure]. tRef. [pretzel1960properties]. uRef. [messer1960survey]. vRef. [zhao2017high]. wRef. [bird2020thermal]. xRef. [martins1990equations]. yRef. [shull1948neutron].
An example of the application of Guyer’s criterion for $L$ = 8.3 mm is shown in Fig. 4. The hydrodynamic window is located up to 15.5 K for PBE, up to 16 K for PBESOL, and up to 17.2 K for LDA, which agrees with the range of temperatures for the experimental observation of second sound in NaF [jackson1970second]. These ranges indicate the temperature intervals within which each functional predicts phonon hydrodynamics. We characterize phonon hydrodynamics windows at characteristic lengths of $L$ = 10 mm, $L$ = 10 $\mu\mathrm{m}$, and $L$ = 10 nm for NaF, LiF, LiH, and NaH (see Appendix D of the supplementary material). We also present the modewise Normal and Umklapp scattering rates at $T$ = 20 K and $T$ = 300 K for the same materials (see Appendix C of the supplementary material).
Figure 4: Phonon hydrodynamics windows for NaF at $L$ = 8.3 mm. Calculations were done with PBE (black dashed lines), PBESOL (red dashed lines), and LDA (blue dashed lines). Double arrow lines show the hydrodynamics range for each functional. The experimental value (Ref. [jackson1970second]) is shown as a dashed vertical orange line.
By applying Guyer’s criterion at different temperatures and characteristic lengths, the ‘phase space’ of phonon transport regimes can be determined. Fig. 5 shows the three phonon transport regimes (ballistic, hydrodynamic, and diffusive) for the eight fluorides, chlorides, and hydrides, over temperatures up to $80\,\mathrm{K}$ and characteristic lengths from $10^{2}$ to $10^{8}\,\mathrm{nm}$. To the best of our knowledge, these are the first reported predictions of phonon hydrodynamics in LiH, NaH, KH, KF, KCl and NaCl. Additionally, we solve the linearized BTE with the full scattering matrix operator for NaF, NaH, LiF, KF and KH to obtain predictions for the observation of second sound (see supplementary material Appendix G). LDA generally yields the widest phonon hydrodynamic window compared to PBE and PBESOL. LiF exhibits the broadest phonon hydrodynamic window, whereas KCl and NaH show the narrowest.
@@
Finally, we briefly comment on the impact of isotopes on the phonon hydrodynamics windows. Resistive scattering rate (R-scattering) can be used instead of Umklapp scattering rate (U-scattering) to predict phonon hydrodynamics by applying Eq. 2. We report the impact of isotopes on the phonon hydrodynamics windows of NaF, LiF, LiH, and NaH (see Fig. 6 and Fig. E.14 in the supplementary material. For fluorides and chlorides, we notice the hydrodynamics window range at $L$ = 1 mm of LiF narrows to 9 K (shaded red region), compared with U-scattering, which extends up to 24.5 K (shaded transparent red region). This is consistent with the analysis from Ackerman and Guyer  [ackerman1968temperature], who argue that high crystal purity is required to observe second sound in LiF. On the other hand, the curves of U-scattering and R-scattering in NaF remain close to each other, and consequently, the hydrodynamics windows are nearly the same. This small impact of isotopes may explain the success of the early experimental work on NaF, confirming the existence of second sound [jackson1970second, mcnelly1970heat]. For hydrides, the hydrodynamic windows disappear at $L$ = 1 mm for LiH when isotope scattering is included, while NaH remains unaffected as NaF. Fig. E.14 also shows the hydrodynamic windows for Si and Ge [huberman2022question] evaluated at the same characteristic length.

### v3

@@ 3.1 General Properties
$Z^{*}$ 0.99 0.98 0.97 1.00 (GGA)ba, 0.98 (LDA)ba
$\omega_{\mathrm{LO}}$ (THz) 22.78 23.35 24.31
$\omega_{\mathrm{TO}}$ (THz) 12.93 14.14 15.67 14.80 (CN)az, 12.60 (QH)az
$E_{g}$ (eV) 3.43 3.25 3.04 5.85 (GW)bc, 3.18 (LDA)bc
aRef. [142]. bRef. [46]. cRef. [101]. dRef. [120]. eRef. [78]. fRef. [99]. gRef. [16]. hRef. [74]. iRef. [113]. jRef. [11]. kRef. [18]. lRef. [143]. mRef. [15]. nRef. [72]. oRef. [71]. pRef. [55]. qRef. [33]. rRef. [144]. sRef. [8]. tRef. [29]. uRef. [42]. vRef. [112]. wRef. [107]. xRef. [98]. yRef. [17]. zRef. [47]. aaRef. [26]. abRef. [90]. acRef. [35]. adRef. [88]. aeRef. [51]. afRef. [130]. agRef. [103]. ahRef. [81]. aiRef. [52]. ajRef. [25]. akRef. [34]. alRef. [138]. amRef. [145]. anRef. [139]. aoRef. [102]. apRef. [79]. aqRef. [5]. arRef. [10]. asRef. [14]. atRef. [68]. auRef. [109]. avRef. [67]. awRef. [100]. axRef. [65]. ayRef. [124]. azRef. [76]. baRef. [7]. bbRef. [58]. bcRef. [69].
@@ 3.2 Mechanical Properties
The mechanical properties of fluorides, chlorides and hydrides, including the bulk modulus and elastic constants, are summarized in table 2. For cubic crystals, the three independent elastic constants are $C_{11}$, $C_{12}$, and $C_{44}$. Our calculated bulk modulus and elastic constant values are in good agreement with the experimental values. We computed the bulk modulus $B$ and the elastic constants $C_{ij}$ at $0\,\mathrm{K}$. LDA typically yields the largest bulk moduli, PBEsol is intermediate, and PBE gives the smallest (e.g., LiF: $B_{\mathrm{PBE}}=67.17\,\mathrm{GPa}$, $B_{\mathrm{PBEsol}}=72.55\,\mathrm{GPa}$, $B_{\mathrm{LDA}}=86.78\,\mathrm{GPa}$) . This ordering follows from their equilibrium volumes: LDA overbinds (smaller lattice constants), which increases the curvature of the energy–volume relation $B_{0}=V_{0}\,\left.\mathrm{d}^{2}E/\mathrm{d}V^{2}\right|_{V_{0}}$ [105]; PBE underbinds (larger volumes), which reduces that curvature, and PBEsol was tuned for solids and therefore falls between. Although $B_{0}$ includes the factor $V_{0}$, the functional trend is governed primarily by the stiffness of the energy–volume curve at equilibrium: LDA overbinding makes the $E(V)$ minimum noticeably narrower (larger curvature), and this increase in curvature more than offsets the smaller $V_{0}$, resulting in systematically larger bulk moduli. Replacing Li with Na and then K systematically softens the lattice at a fixed anion. For fluorides, using PBE for comparisons, the bulk modulus decreases from $B=67.17\,\mathrm{GPa}$ (LiF) to $46.79\,\mathrm{GPa}$ (NaF) and $29.01\,\mathrm{GPa}$ (KF). For chlorides, it decreases from $23.67\,\mathrm{GPa}$ (NaCl) to $16.27\,\mathrm{GPa}$ (KCl). For hydrides, it decreases from $34.24\,\mathrm{GPa}$ (LiH) to $22.16\,\mathrm{GPa}$ (NaH) and $13.74\,\mathrm{GPa}$ (KH). In magnitude, fluorides are the stiffest for a given cation, chlorides are intermediate, and hydrides are the softest for Na and K.
@@
For cubic crystals, mechanical stability requires the elastic constants to satisfy four conditions: $C_{11}>0$, $C_{44}>0$, $C_{11}-C_{12}>0$, and $C_{11}+2C_{12}>0$ [55]. The calculated values of the elastic constants for the fluoride, chloride, and hydride materials satisfy all these conditions. Comparing theoretical elastic constants with experiments is challenging because of the uncertainties in experimentally determining $C_{12}$, which is a linear combination of the elastic constants in the ¡110¿ direction, whereas $C_{11}$ and $C_{44}$ values can be measured directly via ultrasonic techniques[24]. Directional elasticity is captured by the Zener anisotropy factor $AF=2C_{44}/(C_{11}-C_{12})$, where $AF=1$ for an elastic isotropic medium [147]. Using PBE elastic constants, the Zener anisotropy factors are: LiF $AF\!\approx\!1.90$, LiH $AF\!\approx\!1.65$, and NaH $AF\!\approx\!1.45$ are anisotropic; KH $AF\!\approx\!0.94$ is nearly isotropic; NaF $AF\!\approx\!0.72$, NaCl $AF\!\approx\!0.69$, KF $AF\!\approx\!0.61$, and KCl $AF\!\approx\!0.44$ are isotropic. This matches the $AF$ values reported by Sirdeshmukh et al [116].
@@
From the elastic constants, one can also inspect the data to determine whether the material is brittle or ductile. According to Cauchy pressure ($C_{12}$$C_{44}$), if the difference is positive, the material is ductile; otherwise, it is brittle [97]. Based on this, all eight materials are brittle (i.e., negative difference).
@@
Table 2: Bulk modulus ($B$) and elastic constants ($C_{ij}$)of calculated (PBE, PBEsol, LDA), other theoretical data, and experimental data for NaF, LiF, KF, NaCl, KCl, LiH, NaH and KH.
@@
$B$ (GPa) 67.17 72.55 86.78 87.11c, 75.90e, 67.71 (GGA)j, 73.59 (LDA)j, 72.99 (Birch–Murnaghan EOS)k, 70.53 (Vinet EOS)k 66.40f, 76.90g
$C_{11}$ (GPa) 109.45 126.71 143.12 125.80e, 224.83i, 153.15k 113.73f, 135.80g
$C_{12}$ (GPa) 46.03 45.48 50.61 50.90e, 121.09i, 45.62k 47.59f, 47.40g
$C_{44}$ (GPa) 60.32 61.91 68.07 76.10e, 121.09i, 55.08k 63.68f, 68.70g
@@
$B$ (GPa) 13.74 14.67 17.40 13.40 (GGA)p, 17.40 (LDA)p, 13.30q, 19.00r 16.00r, 15.60 $\pm$ 1.50aa
$C_{11}$ (GPa) 28.32 31.81 39.50 39.40p, 26.80q
$C_{12}$ (GPa) 6.45 6.11 6.35 6.30p, 6.50q
$C_{44}$ (GPa) 10.32 10.09 10.43 10.40p, 10.60q
aRef. [142]. bRef. [120]. cRef. [87]. dRef. [104]. eRef. [101]. fRef. [80]. gRef. [48] (data taken from Ref. [101]). hRef. [70] (bulk modulus values from Ref. [120], measured at $T=4.2$ K). iRef. [110]. jRef. [71]. kRef. [55]. lRef. [35]. mRef. [25]. nRef. [150]. oRef. [13]. pRef. [7]. qRef. [149]. rRef. [106]. sRef. [36]. tRef. [89]. uRef. [45]. vRef. [32]. wRef. [125]. xRef. [124]. yRef. [86]. zRef. [31]. aaRef. [53].
@@ 3.3 Thermal Properties
We calculate the room temperature lattice thermal conductivity, specific heat, average Grüneisen parameter, and density values at room temperature. The computed values are listed in table 3 and compared with theoretical and experimental values from the literature. PBEsol generally yields thermal conductivities $k$ that lie between those predicted by PBE (lower) and LDA (higher). This ordering is consistent with LDA’s stronger binding (smaller equilibrium volume), which typically stiffens the phonon spectrum and increases phonon transport (via higher group velocities and longer lifetimes), whereas PBE’s weaker binding shifts $k$ downward, with PBEsol remaining intermediate. Some examples include: for NaF, PBE (18.50 W m-1 K-1) matches experiment (18.40, 18.70) [82, 119], while LDA overestimates (31.50); for KCl, PBEsol (6.90 W m-1K-1) tracks experiments (7.10, 6.70) [82, 116]; for KF and NaCl, LDA (7.32 and 7.86 W m-1 K-1) is the closest to the measured values (7.10) [82, 116]. This variation indicates that the best agreement is material-dependent, because each functional shifts the predicted lattice constant and phonon spectrum differently, which in turn changes both phonon velocities and the strength of anharmonic scattering and, consequently, impacts the thermal conductivity. Specific heats are generally in good agreement with the experimental values across functionals (e.g., for NaCl $c_{p}\approx 0.82\text{--}0.83~\mathrm{kJ\,kg^{-1}\,K^{-1}}$ vs. $0.86~\mathrm{kJ\,kg^{-1}\,K^{-1}}$ [49]). This limited XC sensitivity is expected because $c_{p}$ is governed primarily by the overall vibrational degrees of freedom and is therefore only weakly affected by modest, functional-induced shifts in the phonon spectrum. Average Grüneisen parameters fall within 1.4–2.1 for all materials. On the other hand, density orders across the series (fluorides highest, chlorides intermediate, hydrides lowest), and the computed values closely track the experimental results (PBEsol performs best for halides, while PBE performs best for hydrides). For instance, LiF, PBEsol predicts a value of 2.67 g cm-3, closely matching experiment (2.639) [80], whereas KH, PBE predicts a value of 1.44 g cm-3, closely matching experiments (1.43) [49, 84].
@@
The lattice thermal conductivity is calculated for temperatures ranging from 100 K to 1000 K (see figure  3) and compared with the available data from the literature. The effect of isotopes on the thermal conductivity calculations is also shown in the same figure. All compounds show the expected monotonic decrease of $k$ with $T$. Including naturally abundant isotope scattering in the calculations mainly depresses $k$ (below 200 K) and has less impact at high $T$, where the three functionals produce nearly parallel slopes. Isotope scattering can significantly impact the lattice thermal conductivity $k$ in materials with mixed natural isotopes, but has little to no effect when the constituent elements are effectively monoisotopic. For example, NaF shows essentially no isotope effect because the natural abundance of Na is $\sim 100\%$ $^{23}\mathrm{Na}$ and the natural abundance of F is $\sim 100\%$ $^{19}\mathrm{F}$, whereas LiF exhibits a reduction in $k$ when isotopes are included due to its mixed lithium composition ($\sim 92.5\%$ $^{7}\mathrm{Li}$, $\sim 7.5\%$ $^{6}\mathrm{Li}$ [133]) that introduces phonon mass-disorder scattering most evident at low $T$.
@@
Figure 3: Calculated lattice thermal conductivity for: (a) NaF. (b) LiF. (c) KF. (d) NaCl. (e) KCl. (f) LiH. (g) NaH. (h) KH. Calculations were done with PBE (black solid line), PBEsol (red solid line), and LDA (blue solid line). Calculations also show the effect of the isotopes on the thermal conductivity. Experimental data: Ref. [119] (orange circles), Ref. [96] (gray triangles), Ref. [141] (pink squares), Ref. [95] (orange triangle), Ref. [140] (pink triangles), Ref. [118] (orange squares). Theoretical data: Ref. [37] (purple diamonds), Ref. [72] (pink diamonds), Ref. [71] (green pentagons), Ref. [115] (purple squares), Ref. [145] (purple pentagons).
@@
Table 3: Thermal conductivity ($k$), specific heat ($c_{p}$), average Grüneisen parameter ($\bar{\gamma}$), and density ($\rho$) of calculated (PBE, PBEsol, LDA), other theoretical data, and experimental data for NaF, LiF, KF, NaCl, KCl, LiH, NaH and KH at room temperature.
@@
$k$ ($\mathrm{W\,m^{-1}\,K^{-1}}$) 8.88 11.07 13.91 19.40a, 8.71b, 22.42 ($a_{0}=3.89$ Å)j, 14.59 ($a(300\,\mathrm{K})=4.00$ Å)j, 13.89 ($a_{0}=4.00$ Å)k, 13.48 ($a(300\,\mathrm{K})=4.03$ Å)k 17.60a, 14.09l, 14.20m
$c_{p}$ ($\mathrm{kJ\,kg^{-1}\,K^{-1}}$) 1.62 1.59 1.55 1.62d 1.54d, 1.62m
$\bar{\gamma}$ 2.08 1.88 1.54 2.02b, 1.48e, 1.24 ($a_{0}=4.00$ Å)k, 1.36 ($a(300\,\mathrm{K})=4.03$ Å)k 1.50a, 1.64e
$\rho$ ($\mathrm{g\,cm^{-3}}$) 2.56 2.67 2.88 2.64f 2.64h, 2.64m
@@
$k$ ($\mathrm{W\,m^{-1}\,K^{-1}}$) 4.23 3.31 10.66
$c_{p}$ ($\mathrm{kJ\,kg^{-1}\,K^{-1}}$) 0.99 0.98 0.94
$\bar{\gamma}$ 1.59 1.55 1.28 2.00 (CN)x, 2.20 (QH)x
$\rho$ ($\mathrm{g\,cm^{-3}}$) 1.44 1.51 1.63 1.43q 1.43i, 1.43q
aRef. [82]. bRef. [132]. cRef. [119]. dRef. [134]. eRef. [64]. fRef. [121]. gRef. [70]. hRef. [80]. iRef. [49]. jRef. [72]. kRef. [71]. lRef. [95]. mRef. [116]. nRef. [145]. oRef. [140]. pRef. [146]. qRef. [84]. rRef. [60]. sRef. [36]. tRef. [102]. uRef. [79]. vRef. [152]. wRef. [9]. xRef. [76]. yRef. [114]. $k^{\ast}$ at $T=327$ K.
An example of the application of Guyer’s criterion for $L$ = 8.3 mm is shown in figure 4. The hydrodynamic window is located up to 15.5 K for PBE, up to 16 K for PBEsol, and up to 17.2 K for LDA, which agrees with the range of temperatures for the experimental observation of second sound in NaF [61]. These ranges indicate the temperature intervals within which each functional predicts phonon hydrodynamics. We characterize phonon hydrodynamics windows at characteristic lengths of $L$ = 10 mm, $L$ = 10 $\mu\mathrm{m}$, and $L$ = 10 nm for NaF, LiF, LiH, and NaH (see Appendix E of the supplementary material). We also present the Normal and Umklapp scattering rates at $T$ = 20 K and $T$ = 300 K (see Appendix D of the supplementary material).
Figure 4: Phonon hydrodynamics windows for NaF at $L$ = 8.3 mm. Calculations were done with PBE (black dashed lines), PBEsol (red dashed lines), and LDA (blue dashed lines). Double arrow lines show the hydrodynamics range for each functional. The experimental value (Ref. [61]) is shown as a dashed vertical orange line.
By applying Guyer’s criterion at different temperatures and characteristic lengths, the ‘phase space’ of phonon transport regimes can be determined. figure 5 shows the three phonon transport regimes (ballistic, hydrodynamic, and diffusive) for the eight fluorides, chlorides, and hydrides, over temperatures up to $80\,\mathrm{K}$ and characteristic lengths from $10^{2}$ to $10^{8}\,\mathrm{nm}$. To the best of our knowledge, these are the first reported predictions of phonon hydrodynamics in LiH, NaH, KH, KF, KCl and NaCl. Additionally, we solve the linearized BTE with the full scattering matrix operator for NaF, NaH, LiF, KF and KH to obtain predictions for the observation of second sound (see supplementary material Appendix H). LDA generally yields the widest phonon-hydrodynamic window because, relative to PBE/PBEsol, it shifts the phonon properties in a way that makes Normal scattering dominate over resistive scattering across a broader temperature interval. LiF exhibits the broadest phonon hydrodynamic window, whereas KCl and NaH show the narrowest.
@@
Finally, we briefly comment on the impact of isotopes on the phonon hydrodynamics windows. Resistive scattering rate (R-scattering) can be used instead of Umklapp scattering rate (U-scattering) to predict phonon hydrodynamics by applying Eq. 2. We report the impact of isotopes on the phonon hydrodynamics windows of NaF, LiF, LiH, and NaH (see figure 6 and figure S15 in the supplementary material. For fluorides and chlorides, we notice the hydrodynamics window range at $L$ = 1 mm of LiF narrows to 9 K (shaded red region), compared with U-scattering, which extends up to 24.5 K (shaded transparent red region). This is consistent with the analysis from Ackerman and Guyer  [2], who argue that high crystal purity is required to observe second sound in LiF. On the other hand, the curves of U-scattering and R-scattering in NaF remain close to each other, and consequently, the hydrodynamics windows are nearly the same. This small impact of isotopes may explain the success of the early experimental work on NaF which confirmed the existence of second sound [61, 77]. Likewise, NaH remains unaffected by isotope scattering. On the other hand, the hydrodynamic windows disappear at $L$ = 1 mm for LiH when isotope scattering is included.